Crosshill Companys total overhead costs at various levels of activity are presented below: Month April May Machine- Hours 492. Using the high-low method, estimate a cost formula for maintenance. (Do not round your intermediate calculations. Round th3. Express the companys total overhead cost in the form Y= a +bX. (Do not round your intermediate calculations. Round the V
Answer #1

Answer-1-Compute the maintenance cost in the total overhead cost of $260,130:-

Total Overhead Cost=Utilities +Supervisory salaries +Maintenance

$260,130=$89,700+$70,000+Maintenance

Maintenance= $260,130-$89,700-$70,000

=$100,430

Note: Utilities cost per day= $50,700/ 39,000*69,000

=$89,700

2-Using the high-low method the cost formula for maintenance is:-

Maintenance cost =$19,700( 1.17* Machine Hours)

Working Note:-

The change is high low method is calculated as:-

Machine Hours Maintenance Cost
High activity level 69,000 $100,430
Low activity level 39,000 $65,330
Change 30,000 $35,100

Variable cost per unit= Cost at high activity level-Cost at low activity level/ Units at high activity -Units at low activity

=$35,100/ 30,000

=$1.17 per hour

Fixed cost= Total cost-variable cost element

=100,430-(1.17*69,000)

=100,430-80,730

=$19,700

3-Express the company's total overhead cost in the linear equation form= Y=a+bx

Total overhead cost= Y

Total fixed cost= a

Total Variable overhead per hour= b

Required Machine hours=x

Y=a+bx

Y=$89,700+$2.47 * Machine hours

Working Notes:-

Total Fixed cost= supervisiory salaries+fixed component in mixed maintenance cost

=$70,000+$19,700

=$89,700

Total Variable cost= utilities per hour+ variable component per hour in mixed maintenance cost

=$1.17+$1.3

=$2.47 per machine hour

4-The total factory overhead cost for 44,000 machine hours:-

Machine Hours 44,000
Variable Cost:-
Utilities ($1.3*44,000) $57,200
Fixed Cost:-
Supervisor Salary $70,000
Maintenance Cost:-
Fixed Cost $19,700
Variable Cost ($1.17*44,000) $51,480
Total Cost $198,380

Kindly give me thumbs up if u like my answer...Thanks!!